OpenStudy (anonymous):

For the following equation, log(y) - log(y+1.47) + log13.07 = 2log(x) Express y explicitly as a function of x. Hence, determine the value of y when x = 2.81, giving your answer to 3 decimal places.

jhonyy9 (jhonyy9):

hey ,,yong85" check please your before this exersice --- i have maked these calcules till the end - hope so much that is right and understandably sure good luck bye

OpenStudy (dumbcow):

combine the logs, so it is in the form log(?) = log(?) --> log(y/y+1.47) = log(x^2/13.07) now set the inputs equal y/y+1.47 = x^2/13.07 cross multiply 13.07y = x^2(y) +1.47x^2 13.07y -x^2(y) = 1.47x^2 y(13.07 -x^2) = 1.47x^2 --> y = 1.47x^2/(13.07- x^2) when x=2.81, y = 2.243

OpenStudy (anonymous):

all right thanks! i manage to get it now..